A complex number z satisfies the equation 

z^2\:=3+4i.


Which of the following is true about z?


  1. (3+4i)^2=25+24i
  2. z=+-(2+i)
  3. (3+4i)^2=9+40i
  4. z=+-(2+2i)
  5. (3+4i)^2=7+24i
1
Expert's answer
2021-10-21T10:08:53-0400

Taking into account that

"3+4i=4-1+4i=2^2+4i+i^2=(2+i)^2,"

we conclude that "z^2=(2+i)^2," and hence "z=\pm (2+i)."


Answer: 2